OpenStudy (anonymous):

I need help How can I find the X and y value from the equation? (x - 10^2)+(y-10^2) = 4^2, where (10,10) center of the circle and radius 4.

OpenStudy (anonymous):

I don't think you have the equations quite right should be (x - 10)^2+(y-10)^2 = 4^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i type it wrong

OpenStudy (anonymous):

let't do a simple case u^2+v^2=16 solve for y v^2= 16-u^2 \[v=\pm\sqrt{16-u^2}\]

OpenStudy (anonymous):

u=(x-10) v=y-10 \[(y-10)=\pm\sqrt{16-(x-10)^2}\] \[(y)=\pm\sqrt{16-(x-10)^2}+10\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

that's what you needed right?

OpenStudy (anonymous):

actually I need to find a point (X,Y) above the circle

OpenStudy (anonymous):

I think I need to use iteration for x value to find y